# Rapid, strongly magnetized accretion in the zero-net-vertical-flux shearing box

### v1

@@ 6 Sustaining the low-$\beta$ state
In this section we study how the low-$\beta$ state is maintained via a dynamo feedback that regenerates flux lost through the boundaries. We do not considwe the high-$\beta$ cycles, since this has been studied extensively in past work (e.g., Lesur & Ogilvie, 2008b; Guan & Gammie, 2011; Squire & Bhattacharjee, 2015a; Gressel & Pessah, 2015), focusing instead on how the system can maintain strong time-independent mean fields. (Note that the term “dynamo” is sometimes used to refer specifically to cyclic behavior in the accretion-disk literature, but here we use it more generally to describe the mechanism by which a large-scale field is generated.) While we provide clear evidence that fields are regenerated and expelled on several-orbit time-scales — viz., the system does not reach steady state because its activity shuts down, but because flux loss and growth are balanced — the general principles of this mechanism’s operations, and how this depends on parameters, remain poorly understood.
@@ 6.1 Mean-field dynamo: background
Figure 12: Terms in the equation for $\partial_{t}\overline{B_{y}}$ in $\beta$R0.1, averaged over the steady state in time (see Eq. 10). We see that near the midplane, $B_{y}$ flux is supplied by the shearing of $\overline{B_{x}}$, which is balanced by transport to higher $|z|$ from the turbulence ($\partial_{z}\mathcal{E}_{x}$) with small contributions from mean outflows on either side of the midplane. In contrast, at large $|z|$, $\overline{B_{y}}$ is sustained primarily by the upwards transport of flux by the turbulence ($\partial_{z}\mathcal{E}_{x}$), which competes against strong vertical outflows through the boundary. This shows that, absent a dynamo effect to regenerate $\overline{B_{x}}$, the field would decay within several orbits. The thin purple line illustrates the good comparison between measured $\partial_{z}\mathcal{E}_{x}$ and its expectation based on standard dynamo theory $\mathcal{E}_{x}\approx\eta_{\rm turb}\partial_{z}B_{y}$ with $\eta_{\rm turb}=\Omega^{-1}\overline{\delta\bm{u}^{2}}/3$ (we smooth this with a 5-gridpoint moving average for clarity).
@@
where the ‘transport coefficients’ $\alpha_{D}$ and $\eta_{D}$ are determined by the dependence of the fluctuations on the mean fields. The second expression results from assuming homogeneous and isotropic fluctuations, which, while not valid in a stratified shear flow, is helpful for the simple analysis carried out here ($\alpha_{D}$ is the diagonal part of $\alpha_{D,ij}$; $\eta_{D}$ is the fully antisymmetric part of $\tilde{\eta}_{ijk}$). The so-called $\alpha$-$\Omega$ dynamo results for non-zero $\alpha_{D}$, when the strong $\overline{B_{y}}$ generates $\overline{B_{x}}$ via the final term in Eq. 10a, which then regenerates $\overline{B_{y}}$ via $-q\Omega\overline{B_{x}}$. The effect can be understood as a linear instability by ignoring the effect of $\overline{U_{z}}$ and inserting Eq. 11 into (10), then assuming $\overline{B_{i}}\propto e^{i\kappa z}$ and $|\alpha_{D}\kappa|\ll q\Omega$; one finds solutions with frequency $\omega_{D}\approx-i\kappa^{2}\eta_{D}\pm\sqrt{i\kappa\alpha_{D}q\Omega}$, such that a branch of growing and oscillating solutions exists at sufficiently small $\kappa$, where the dynamo feedback (via $\alpha_{D}$) overcomes the turbulent diffusion (via $\eta_{D}$). The crux, of course, is having fluctuations that cause a non-zero $\alpha_{D}$ of the correct sign, which can result from the combination of stratification and rotation creating a net helicity $\overline{\delta\bm{u}\cdot\nabla\times\delta\bm{u}}$ or current helicity $\overline{\delta\bm{B}\cdot\nabla\times\delta\bm{B}}$. This formalism, or extensions of it, has been analysed in a number of papers in order to understand the cyclic behavior of the high-$\beta$ state (see, e.g., Gressel & Pessah 2015; Tharakkal et al. 2023 and references therein). It also possible to have a similar large-scale dynamo with $\alpha_{D}=0$ as a result of off-diagonal components of $\tilde{\eta}_{D,ijk}$ (Rädler & Stepanov, 2006; Squire & Bhattacharjee, 2015b).
@@ 6.2 Sustaining a strong azimuthal field against escape
Mean-field theory is applied to the saturated low-$\beta$ state in Fig. 12. We illustrate the time average over the steady state of each term in Eq. 10b, normalizing each by $\overline{B_{y}}(z)/2\pi$, such that a value of $1$ implies that the particular term would grow or decay $\overline{B_{y}}$ by a factor $e^{\pm 1}$ over one orbit at a given $z$. The sum over all terms is effectively zero, as must be the case in steady state.
Our general conclusions are similar to Johansen & Levin (2008) (figures 11-14), with the important difference that flux is continuously escaping the boundaries in our case. Fields are replenished and expelled over several-orbit timescales, with strong flux creation by shearing for $H_{\rm th}\lesssim|z|\lesssim 3$ balanced by its diffusion to larger radii by turbulence ($\partial_{z}\mathcal{E}_{x}$). For $|z|\gtrsim 3$, $\overline{B_{y}}$ is instead replenished by fluctuation-induced diffusion from smaller $z$ (i.e., turbulent transport; $\partial_{z}\mathcal{E}_{x}$), while being rapidly carried out of the box by the expanding outflows ($\overline{B_{y}}\partial_{z}\overline{U_{z}}$; advection by $\overline{U_{z}}$ is only a minor effect because $\overline{U_{z}}$ is purely compressive). With the thin purple line, we compare the measured $\partial_{z}\mathcal{E}_{x}$ from $\overline{\delta\bm{u}\times\delta\bm{B}}$ to the expectation for $\eta_{D}$ from quasi-linear dynamo theory, $\eta_{D}=\overline{\delta\bm{u}^{2}}/3\Omega$, assuming the correlation time of the turbulence is ${\simeq}\Omega^{-1}$ (there is no contribution to $\eta_{D}$ from $\delta\bm{B}$ fluctuations in quasi-linear theory; Rincon, 2019). We see surprisingly good agreement, with $\partial_{z}\mathcal{E}_{x}\approx\partial_{z}(\eta_{D}\partial_{z}\overline{B_{y}})$ across the full domain; this indicates that the fluctuations are predominantly affecting $\overline{B_{y}}$ as a turbulent diffusion, as expected (the effect of a possible $\alpha_{D}$ on $\partial_{t}\overline{B_{y}}$ would be very small, being proportional to $\overline{B_{x}}$).
@@
Figure 13 shows the same $\overline{B_{y}}$-dynamo analysis for two other cases of relevance. The first, $\beta$0.1-H4, is the case discussed in §5.6 that starts from $\beta\ll 1$ and loses all of its flux until reaching $\beta\sim 1$ in the midplane. We consider the middle of this flux-decrease phase, averaging from $t\Omega/2\pi\approx 6$ to 8 (fourth to sixth line from the top in Fig. 10a), but the results are similar at all times during this phase. As expected, the total $\overline{B_{y}}$ contribution from all terms is nearly constant and negative, indicating $\overline{B_{y}}$ is decreasing in time without changing shape. But, we also still see robust field creation for $|z|\gtrsim H_{\rm th}$, and the balance of terms is similar to the true saturated state in Fig. 12. In other words, during this flux-decrease phase, a similar dynamo continues to operate, but the growth via $q\Omega\overline{B_{x}}$ is overwhelmed by the turbulent diffusion and outflows. The properties for $|z|\lesssim H_{\rm th}$ are explained by our previous observations that the midplane density is increasing by inwards collapse (thereby requiring a large positive contribution from $\overline{U_{z}}$) but the midplane $\overline{B_{y}}$ is not (thereby requiring a large negative contribution from $\overline{B_{x}}$ to balance that from $\overline{U_{z}}$). How the system conspires to reverse the midplane $\overline{B_{x}}$ in order to achieve this is not at all obvious, but it is nonetheless effectively guaranteed if $\overline{B_{y}(z,t)}$ is to maintain a similar spatial form as the total flux decreases.
@@
The second example in Fig. 13 is $\beta$0.001, which maintains a $\beta\ll 1$ midplane because the collapse of density is halted by the simulation’s low resolution per $H_{\rm th}$. As discussed in §5.6, while it is certainly not surprising that the grid can stop the density’s collapse to ${\simeq}H_{\rm th}$, the fact that this in turn stops the escape of flux is more interesting. Indeed, we see that its dynamo continues to operate, supporting the field against the turbulent diffusion in the midplane and rapid outflows to the boundaries. While the loss/growth rate is somewhat lower than in the resolved case ($\beta$R0.1; Fig. 12), it still implies complete regeneration of the flux over $4$ to $5$ orbits in regions away for the midplane that should, in principle, be ignorant of $H_{\rm th}$ and thus adequately resolved. Note also that the near-midplane turbulence in this simulation is highly supersonic, as needed to cause a similar relative $\partial_{z}\mathcal{E}_{x}$ at $\beta\ll 1$.
@@
A simple interpretation of these results is that, in principle, any $\beta\ll 1$ state can be maintained, with the dynamo generating a $\overline{B_{x}}$ near the midplane that in turn balances the continual loss of $\overline{B_{y}}$ through the boundaries. However, in the absence of a reason for the density to choose a particular scale height $H_{\rho}$, it always collapses over orbital timescales towards a thermal profile in the midplane, with the magnetic field (and thus the dynamo) being forced to cooperate. It stops once $P_{\rm th}$ provides significant pressure support, corresponding to the $\beta\sim 1$ midplane we see in the steady state of well-resolved simulations. On the other hand, if other effects intervene — the grid in our simulations, but perhaps physical midplane turbulence of another source in more realistic settings — the collapse is halted with $H_{\rho}\gg H_{\rm th}$, allowing a $\beta\ll 1$ midplane.
@@ 6.3 Generation of the radial field
The results of Eqs. 10b and 13 show only that $\overline{B_{y}}$ is driven primarily by $\overline{B_{x}}$, providing no information on $\partial_{t}\overline{B_{x}}$, which must also be continuously regenerated to balance its escape. Unfortunately a similar analysis for $\overline{B_{x}}$ is less useful: $-\partial_{z}\mathcal{E}_{y}$ can provide both turbulent diffusion through $\eta_{D}$, and growth of $\overline{B_{x}}$ through $\mathcal{E}_{y}\sim\alpha_{D}\overline{B_{y}}$; absent other contributions from mean flows, the two effects will generally cancel so that $\partial_{t}\overline{B_{x}}\approx 0$, implying that $\partial_{z}\mathcal{E}_{y}\approx 0$ provides no information about the relative sizes of growth and diffusion terms. While a number of techniques have been developed to overcome this and measure $\alpha_{D}$ and $\eta_{D}$ in simulations (e.g., the “test-field” method; Schrinner et al. 2005), all have limitations and are rather complex, falling beyond the scope of this paper. Instead, motivated by the apparent success of quasi-linear dynamo theory for predicting $\partial_{z}\mathcal{E}_{x}$ (thin purple line in Fig. 12), here we attempt to test the basic $\alpha$-$\Omega$ dynamo paradigm in which $\alpha_{D}$ is driven by the fluctuations’ helicity, using the simplest, standard quasi-linear result for helical velocity and magnetic fluctuations:
@@
$\alpha_{D}\approx-\frac{1}{3\Omega}(\overline{\delta\bm{u}\cdot\nabla\times\delta\bm{u}}-\overline{\delta\bm{B}\cdot\nabla\times\delta\bm{B}}/\rho),$ (12)
@@
where again we assume correlation time $\Omega^{-1}$ (Pouquet et al., 1976; Rincon, 2019). This $\alpha$ effect then enters the mean induction equation (10a) as $-\partial_{z}(\alpha_{D}\overline{B_{y}})$, so this form of dynamo feedback would require $\alpha_{D}<0$ at $z>0$ since $\overline{B_{x}}<0$ is needed to drive $\overline{B_{y}}>0$ (with $\partial_{z}\overline{B_{y}}<0$; likewise, $\alpha_{D}>0$ at $z<0$). [^9] In principle, the expression $-\partial_{z}(\alpha_{D}\overline{B_{y}})$ could allow a dynamo driven by an $\alpha_{D}$ gradient instead, in which case $\partial_{z}\alpha_{D}>0$ would be needed. However, this is not the standard $\alpha$-$\Omega$ mechanism and many other terms beyond the simple scalar expansion for $\alpha_{D}$ in Eq. 11 may become important also. To estimate the value of $\alpha_{D}$ that would be required to drive the observed dynamo, we can use the approximate power-law equilibrium (4) with $v_{\rm A}/\Omega=z\sqrt{2/a}$ to write $\partial_{z}\mathcal{E}_{y}\sim\alpha_{D}\partial_{z}B_{y}\sim\sqrt{a/2}\Omega\overline{B_{y}}\alpha_{D}/v_{\rm A}$ (where $a/2=|\partial\ln\overline{B_{y}}/\partial\ln z|$). Thus, to regenerate $B_{x}$ in one orbit requires $\alpha_{D}/v_{\rm A}\simeq(2\pi)^{-1}\sqrt{2/a}\,\overline{B_{x}}/\overline{B_{y}}$; we have $|\overline{B_{x}}/\overline{B_{y}}|\lesssim 0.05$ in ZNVF flux cases, suggesting $\alpha_{D}/v_{\rm A}\sim 0.01$ should be sufficient to sustain the dynamo.
@@
We test these ideas in Fig. 14 for the simulations shown in Figs. 12 and 13, all of which exhibit a continuously regenerated $\overline{B_{x}}$. As well as Eq. 12, we plot the contribution from each direction of $\delta\bm{u}$ and $\delta\bm{B}$ fluctuations ($\alpha_{D,u,x}=-\overline{\delta u_{x}(\nabla\times\bm{u})_{x}}/3\Omega$, $\alpha_{D,B,y}=\overline{\delta B_{y}(\nabla\times\bm{B})_{y}}/3\Omega$ etc.) to attempt to extract general features and trends. Most importantly, although the magnitude of $\alpha_{D}$ is in principle sufficiently large to produce the observed dynamo feedback, we see no consistent sign of $\alpha_{D}$ between each of the 3 cases shown, including for $\alpha_{D,u}$ and $\alpha_{D,B}$ individually (this could be relevant, for example, if the magnetic contribution to $\alpha_{D}$ in Eq. 12 was more or less efficient at driving a dynamo than the velocity contribution; see e.g., Gressel 2010; Rincon 2019). Instead, we find empirically that the sign of $\alpha_{D}$ seems to be well predicted by whether $\beta>0.03$ (grey-shaded regions; $\alpha_{D}>0$), or $\beta<0.03$ ($\alpha_{D}<0$), a feature that seems robust across all of our simulations, including those with net vertical flux (not shown). This change is driven primarily by the stronger magnetic fluctuations at low $\beta$, and may relate to a greater dominance of the Parker instability compared to the MRI, although this is hard to diagnose. Of particular relevance is $\beta$R0.1, which, with $\alpha_{D}>0$ at $z>0$, would not sustain the observed $\overline{B_{x}}$ if Eq. 12 were correct, despite apparently having a rather similar dynamo to $\beta$0.1-H4 and $\beta$0.001 (for which $\alpha_{D}<0$ at $z>0$).
@@
We thus conclude that the dynamo mechanism regenerating $\bar{B}_{x}$ is likely not directly related to the standard $\alpha$ effect driven by flow or current helicity. This is consistent with previous simulations of buoyancy-driven large-scale dynamos, which have seen similar inconsistencies between the measured sign of helicity and theoretical expectations (e.g., Tharakkal et al., 2023, and references therein). One possible explanation is that there exist two mechanisms — driven, e.g., by MRI or the Parker instability — which can both provide the required dynamo feedback but in different ways (e.g., in low-$\beta$ Parker-dominated regions, $\alpha_{D}$ does have the required sign). This resembles ideas proposed in Johansen & Levin (2008) as well as previous work on the high-$\beta$ dynamo cycles (Gressel, 2010). Another non-exclusive possibility is that the feedback is unrelated to the helicity, instead relying only on the shear flow via off-diagonal $\tilde{\eta}_{D,ijk}$ terms (Rädler & Stepanov, 2006; Rogachevskii & Kleeorin, 2008; Squire & Bhattacharjee, 2016). This would more easily explain the apparent robustness of the field regeneration, but, other than at low Reynolds numbers (Squire & Bhattacharjee, 2015c), the properties of this type of dynamo remain poorly understood (Käpylä et al., 2020; Zhou & Blackman, 2021).

### v3

@@ 6 Sustaining the low-$\beta$ state
In this section we study how the low-$\beta$ state is maintained via a dynamo feedback that regenerates flux lost through the boundaries. We do not consider the high-$\beta$ cycles, since this has been studied extensively in past work (e.g., Lesur & Ogilvie, 2008b; Squire & Bhattacharjee, 2015a; Gressel & Pessah, 2015), focusing instead on how the system can maintain strong time-independent mean fields. (Note that the term “dynamo” is sometimes used to refer specifically to cyclic behavior in the accretion-disk literature, but here we use it more generally to describe the mechanism by which a large-scale field is generated.) While we provide clear evidence that fields are regenerated and expelled on several-orbit time-scales — viz., the system does not reach steady state because its activity shuts down, but because flux loss and growth are balanced — the physics of this mechanism, and how this depends on parameters, remain poorly understood.
@@ 6.1 Mean-field dynamo: background
where the ‘transport coefficients’ $\alpha_{D}$ and $\eta_{D}$ are determined by the dependence of the fluctuations on the mean fields. The second expression results from assuming homogeneous and isotropic fluctuations, which, while not valid in a stratified shear flow, is helpful for the simple analysis carried out here ($\alpha_{D}$ is the diagonal part of $\alpha_{D,ij}$; $\eta_{D}$ is the fully antisymmetric part of $\tilde{\eta}_{ijk}$). The so-called $\alpha$-$\Omega$ dynamo results for non-zero $\alpha_{D}$, when the strong $\overline{B_{y}}$ generates $\overline{B_{x}}$ via the final term in Eq. 10a, which then regenerates $\overline{B_{y}}$ via $-q\Omega\overline{B_{x}}$. The effect can be understood as a linear instability by ignoring the effect of $\overline{U_{z}}$ and inserting Eq. 11 into (10), then assuming $\overline{B_{i}}\propto e^{i\kappa z}$ and $|\alpha_{D}\kappa|\ll q\Omega$; one finds solutions with frequency $\omega_{D}\approx-i\kappa^{2}\eta_{D}\pm\sqrt{i\kappa\alpha_{D}q\Omega}$, such that a branch of growing and oscillating solutions exists at small $\kappa$, where the dynamo feedback (via $\alpha_{D}$) overcomes the turbulent diffusion (via $\eta_{D}$). The crux, of course, is having fluctuations that cause a sufficiently large $\alpha_{D}$, which can result from the combination of stratification and rotation creating a net helicity $\overline{\delta\bm{u}\cdot\nabla\times\delta\bm{u}}$ or current helicity $\overline{\delta\bm{B}\cdot\nabla\times\delta\bm{B}}$. This formalism, or extensions of it, has been analysed in a number of papers in order to understand the cyclic behavior of the high-$\beta$ state (see, e.g., Gressel & Pessah 2015; Mondal & Bhat 2023 and references therein). It is also possible to have a similar large-scale dynamo with $\alpha_{D}=0$ as a result of off-diagonal components of $\tilde{\eta}_{D,ijk}$ (Rädler & Stepanov, 2006; Squire & Bhattacharjee, 2015b; Mondal & Bhat, 2023).
@@ 6.2 Sustaining a strong azimuthal field against escape
Mean-field theory is applied to the saturated low-$\beta$ state in Fig. 12. We illustrate the time average over the steady state of each term in Eq. 10b, normalizing each by $\overline{B_{y}}(z)/2\pi$, such that a value of $1$ implies that the particular term would grow or decay $\overline{B_{y}}$ by a factor $e^{\pm 1}$ over one orbit at each $z$. The sum over all terms is effectively zero, as must be the case in steady state.
Our general conclusions are similar to Johansen & Levin (2008) (figures 11-14), with the important difference that flux is continuously escaping the boundaries in our case. Fields are replenished and expelled over several-orbit timescales, with strong flux creation by shearing for $H_{\rm th}\lesssim|z|\lesssim 3$ balanced by its diffusion to larger $|z|$ by turbulence ($\partial_{z}\mathcal{E}_{x}$). For $|z|\gtrsim 3$, $\overline{B_{y}}$ is instead replenished by fluctuation-induced diffusion from smaller $z$ (i.e., turbulent transport; $\partial_{z}\mathcal{E}_{x}$), while being rapidly carried out of the box by the expanding outflows ($\overline{B_{y}}\partial_{z}\overline{U_{z}}$; advection by $\overline{U_{z}}$ is only a minor effect because $\overline{U_{z}}$ is purely compressive). With the thin purple line, we compare the measured $\partial_{z}\mathcal{E}_{x}$ from $\overline{\delta\bm{u}\times\delta\bm{B}}$ to the expectated $\eta_{D}$ from quasi-linear dynamo theory, $\eta_{D}=\overline{\delta\bm{u}^{2}}/3\Omega$, assuming the correlation time of the turbulence is ${\simeq}\Omega^{-1}$ (there is no contribution to $\eta_{D}$ from $\delta\bm{B}$ fluctuations in quasi-linear theory; Rincon, 2019). We see surprisingly good agreement, with $\partial_{z}\mathcal{E}_{x}\approx\partial_{z}(\eta_{D}\partial_{z}\overline{B_{y}})$ across the full domain; this indicates that the fluctuations are predominantly affecting $\overline{B_{y}}$ as a turbulent diffusion, as expected (the effect of a possible $\alpha_{D}$ on $\partial_{t}\overline{B_{y}}$ would be very small, being proportional to $\overline{B_{x}}$).
@@
Figure 13 shows the same $\overline{B_{y}}$-dynamo analysis for two other cases of relevance. The first, $\beta$0.1-H4, is discussed in §5.5; it starts from $\beta\ll 1$ and loses all of its flux until reaching $\beta\sim 1$ in the midplane. We consider the middle of this flux-decrease phase, averaging from $t\Omega/2\pi\approx 6$ to 8 (fourth to sixth line from the top in Fig. 10a), but the results are similar at all times during this phase. As expected, the total $\overline{B_{y}}$ contribution from all terms is nearly constant and negative, indicating $\overline{B_{y}}$ is decreasing in time without changing shape. But, we also still see robust field creation for $|z|\gtrsim H_{\rm th}$, and the balance of terms is similar to the true saturated state in Fig. 12. In other words, during this flux-decrease phase, a similar dynamo continues to operate, but the growth via $q\Omega\overline{B_{x}}$ is overwhelmed by the turbulent diffusion and outflows. The properties for $|z|\lesssim H_{\rm th}$ are explained by our previous observations that the midplane density is increasing by inwards collapse (thereby requiring a large positive contribution from $\overline{U_{z}}$) but the midplane $\overline{B_{y}}$ is not (thereby requiring a large negative contribution from $\overline{B_{x}}$ to balance that from $\overline{U_{z}}$). How the system conspires to reverse the midplane $\overline{B_{x}}$ in order to achieve this is not at all obvious, but it is nonetheless effectively guaranteed if $\overline{B_{y}(z,t)}$ is to maintain a similar spatial form as the total flux decreases.
@@
The second example in Fig. 13 is $\beta$0.001, which maintains a $\beta\ll 1$ midplane because the collapse of density is halted by the simulation’s low number of grid cells per $H_{\rm th}$. While it is not surprising that the coarse grid stops the density’s collapse to ${\simeq}H_{\rm th}$, the observation that this also stops the escape of flux is interesting (see also §5.5). Indeed, we see that the system’s dynamo continues to operate, supporting the field against the turbulent diffusion in the midplane and rapid outflows to the boundaries. While the loss/growth rate is somewhat lower than in the resolved case ($\beta$R0.1; Fig. 12), it still implies complete regeneration of the flux over $4$ to $5$ orbits in regions away from the midplane that should, in principle, be ignorant of $H_{\rm th}$ and thus adequately resolved. Note also that the near-midplane turbulence in this simulation is highly supersonic, as needed to cause a similar relative $\partial_{z}\mathcal{E}_{x}$ at $\beta\ll 1$.
@@
A simple interpretation of these results is that, in principle, any $\beta\ll 1$ state can be maintained, with the dynamo generating a $\overline{B_{x}}$ near the midplane that in turn balances the continual loss of $\overline{B_{y}}$ through the boundaries. However, without a reason for the density to stop at a particular scale height $H_{\rho}$, it always collapses towards a thermal profile in the midplane, with the magnetic field (and thus the dynamo) being forced to cooperate. It stops once $P_{\rm th}$ provides significant pressure support, corresponding to the $\beta\sim 1$ midplane in the steady state of well-resolved simulations. On the other hand, if other effects intervene — the grid in our simulations, but perhaps e.g., radiation, or physical midplane turbulence from another source in more realistic settings — the collapse is halted with $H_{\rho}\gg H_{\rm th}$, thereby also enabling a $\beta\ll 1$ midplane.
@@ 6.3 Generation of the radial field
The results of Eqs. 10b and 13 show only that $\overline{B_{y}}$ is driven primarily by $\overline{B_{x}}$, providing no information on $\partial_{t}\overline{B_{x}}$, which must also be continuously regenerated to balance its escape. Unfortunately a similar analysis for $\overline{B_{x}}$ is less useful: $-\partial_{z}\mathcal{E}_{y}$ can provide both turbulent diffusion through $\eta_{D}$, and growth of $\overline{B_{x}}$ through $\mathcal{E}_{y}\sim\alpha_{D}\overline{B_{y}}$; absent other contributions from mean flows, the two effects will generally cancel so that $\partial_{t}\overline{B_{x}}\approx 0$. This implies that $\partial_{z}\mathcal{E}_{y}\approx 0$ does not yield useful information about the relative sizes of growth and diffusion terms. While a number of techniques have been developed to overcome this and measure $\alpha_{D}$ and $\eta_{D}$ in simulations (e.g., the “test-field” method; Schrinner et al. 2005), all have limitations and are rather complex, falling beyond the scope of this paper. Instead, motivated by the apparent success of quasi-linear dynamo theory for predicting $\partial_{z}\mathcal{E}_{x}$ (thin purple line in Fig. 12), here we attempt to test the basic $\alpha$-$\Omega$ dynamo paradigm in which $\alpha_{D}$ is driven by the fluctuations’ helicity, using the simplest, standard quasi-linear result for helical velocity and magnetic fluctuations:
@@
$\alpha_{D}\approx-\frac{1}{3\Omega}(\overline{\delta\bm{u}\cdot\nabla\times\delta\bm{u}}-\overline{\delta\bm{B}\cdot\nabla\times\delta\bm{B}}/\rho).$ (12)
@@
We again assume a correlation time $\Omega^{-1}$ (Pouquet et al., 1976; Rincon, 2019). This $\alpha$ effect then enters the mean induction equation (10a) as $-\partial_{z}(\alpha_{D}\overline{B_{y}})$, so this form of dynamo feedback would require $\alpha_{D}<0$ at $z>0$ since $\overline{B_{x}}<0$ is needed to drive $\overline{B_{y}}>0$ when $\partial_{z}\overline{B_{y}}<0$ (likewise, $\alpha_{D}>0$ at $z<0$). [^11] In principle, the expression $-\partial_{z}(\alpha_{D}\overline{B_{y}})$ could allow a dynamo driven by an $\alpha_{D}$ gradient instead, in which case $\partial_{z}\alpha_{D}>0$ would be needed. However, this is not the standard $\alpha$-$\Omega$ mechanism and many other terms beyond the simple scalar expansion for $\alpha_{D}$ in Eq. 11 may become important also. To estimate the value of $\alpha_{D}$ that would be required to drive the observed dynamo, we can use the approximate power-law equilibrium (4) with $v_{\rm A}/\Omega=z\sqrt{2/a}$ to write $\partial_{z}\mathcal{E}_{y}\sim\alpha_{D}\partial_{z}B_{y}\sim\sqrt{a/2}\Omega\overline{B_{y}}\alpha_{D}/v_{\rm A}$ (where $a/2=|\partial\ln\overline{B_{y}}/\partial\ln z|$). Thus, to regenerate $B_{x}$ in one orbit requires $\alpha_{D}/v_{\rm A}\simeq(2\pi)^{-1}\sqrt{2/a}\,\overline{B_{x}}/\overline{B_{y}}$; we have $|\overline{B_{x}}/\overline{B_{y}}|\lesssim 0.05$ in ZNVF flux cases, suggesting $\alpha_{D}/v_{\rm A}\sim 0.01$ should be sufficient to sustain the dynamo.
@@
We test these ideas in Fig. 14 for the simulations shown in Figs. 12 and 13, all of which exhibit a continuously regenerated $\overline{B_{x}}$. As well as Eq. 12, we plot the contribution from each direction of $\delta\bm{u}$ and $\delta\bm{B}$ fluctuations ($\alpha_{D,u,x}=-\overline{\delta u_{x}(\nabla\times\bm{u})_{x}}/3\Omega$, $\alpha_{D,B,y}=\overline{\delta B_{y}(\nabla\times\bm{B})_{y}}/3\Omega$ etc.) to attempt to extract general features and trends. Most importantly, although the magnitude of $\alpha_{D}$ is in principle sufficiently large to produce the observed dynamo feedback, we see no consistent sign of $\alpha_{D}$ between each of the 3 cases shown, including for $\alpha_{D,u}$ and $\alpha_{D,B}$ individually (this could be relevant, for example, if the magnetic contribution to $\alpha_{D}$ in Eq. 12 was more or less efficient at driving a dynamo than the velocity contribution; see e.g., Gressel 2010; Rincon 2019). Instead, we find empirically that the sign of $\alpha_{D}$ seems to be well predicted by whether $\beta>0.03$ (grey-shaded regions; $\alpha_{D}>0$), or $\beta<0.03$ ($\alpha_{D}<0$), a feature that seems robust across all of our simulations, including those with net vertical flux (not shown). This change is driven primarily by the stronger magnetic fluctuations at low $\beta$, and may relate to a greater dominance of the Parker instability compared to the MRI, although this is hard to diagnose. $\beta$R0.1 is of particular relevance to our overall conclusion: with $\alpha_{D}>0$ at $z>0$, it would not sustain the observed $\overline{B_{x}}$ if Eq. 12 were correct, despite apparently having a rather similar dynamo to $\beta$0.1-H4 and $\beta$0.001 (for which $\alpha_{D}<0$ at $z>0$).
@@
We thus conclude that the dynamo mechanism regenerating $\bar{B}_{x}$ is likely not the standard $\alpha$ effect driven by flow or current helicity. This is consistent with previous simulations of buoyancy-driven large-scale dynamos, which have seen similar inconsistencies between the measured sign of helicity and theoretical expectations (e.g., Tharakkal et al., 2023, and references therein). One possible explanation is that there exist two mechanisms — driven, e.g., by MRI or the Parker instability — which can both provide the required dynamo feedback but in different ways (e.g., in low-$\beta$ Parker-dominated regions, $\alpha_{D}$ does have the required sign). This resembles ideas proposed in Johansen & Levin (2008) and is also consistent with analyses of differences between midplane and coronal regions in the high-$\beta$ state (Gressel, 2010; Held et al., 2024). Similarly, the analysis of Held et al. (2024) suggests that local vertical fields play a key role in low-$\beta$ regions, generating radial fields via vertical gradients in the flow; this interplay would not show up in our analysis to our horizontal average but would be interesting to study further in the low-$\beta$ state. Another non-exclusive possibility is that the feedback is unrelated to the helicity, instead relying only on the shear flow via off-diagonal $\tilde{\eta}_{D,ijk}$ terms (Rädler & Stepanov, 2006; Rogachevskii & Kleeorin, 2008; Squire & Bhattacharjee, 2016). This would more easily explain the apparent robustness of the field regeneration, but, other than at low Reynolds numbers (Squire & Bhattacharjee, 2015c), the properties of this type of dynamo remain poorly understood (Käpylä et al., 2020; Zhou & Blackman, 2021; Mondal & Bhat, 2023).